OpenStudy (anonymous):

A biologist took a count of the number of fish in a particular lake, and recounted the lake’s population of fish on each of the next six weeks. Find a quadratic function that models the data as a function of x, the number of weeks. Use the model to estimate the number of fish at the lake on week 8. P(x) = 13x^2 – 10x + 350; 917 fish P(x) = 13x^2 – 10x + 350; 1,102 fish P(x) = 18x^2 + 10x + 300; 1,252 fish P(x) = 18x^2 + 10x + 300; 1,532 fish

OpenStudy (anonymous):

Week 0 1 2 3 4 5 6

OpenStudy (anonymous):

Population 350 353 382 437 518 625 758

OpenStudy (anonymous):

first, in the given options what is the population in week 0 in each of them ?

OpenStudy (anonymous):

Wouldnt I write it like this : 13x^2 – 10x + 350; 13(0)^2-10(350)+350 ?

OpenStudy (anonymous):

for the first : 13x^2 – 10x + 350 plugging x = 0 13 * (0^2) - 10 * 0 + 350 = .. ?

OpenStudy (anonymous):

You lost me . . . :/

OpenStudy (anonymous):

in order to get from each equation the amount of fish every week we plug x = number of week so in order to get the number of fish in week number 0 we just replace x by 0 in each equation the ones that will be good for us will be the ones with 350 fish in week number 0.

OpenStudy (anonymous):

Oh , i see that . . .

OpenStudy (anonymous):

so what can you tell about the number of fish in week number 0 according to each one of the equations ?

OpenStudy (anonymous):

I dont understand what you are asking ?

OpenStudy (anonymous):

now in order to see which of the equations really shows 350 fish in week 0 we need to plug x =0 in each of them so now we have to check it

OpenStudy (anonymous):

plug 0 into what equation though ?

OpenStudy (anonymous):

into each one of them so we know which of them shows 350 fish in week number 0..

OpenStudy (anonymous):

so it would look like : 13(0)^2 – 10(0) + 350 ?

OpenStudy (anonymous):

yes so what is left ?

OpenStudy (anonymous):

&& then the second one would look like : 13(2)^2 – 10(2) + 350

OpenStudy (anonymous):

it would be just 350 ?

OpenStudy (anonymous):

yes!.. so in the first we have 350 what in the second ? third ? last ?

OpenStudy (anonymous):

the second i got 302 ?

OpenStudy (anonymous):

the second looks exactly as the first so i dont understand how you got 302

OpenStudy (anonymous):

am i supposed to plug in 0 for the second one too ?

OpenStudy (anonymous):

yes.. no we are checking which of the equations shows 350 fish for week number 0 so we will be able to eliminate the ones who shows different number

OpenStudy (anonymous):

do you understand the logic behind this process ? in the given data we had 350 fish in week number 0 so our equation should show this thing as well !

OpenStudy (anonymous):

because we are looking for an equation that represents the data

OpenStudy (anonymous):

I kinda understand it .

OpenStudy (anonymous):

so the 2nd would be: 13(0)^2 – 10(0) + 350

OpenStudy (anonymous):

its the same exact thing as the 1st ?

OpenStudy (anonymous):

look at the equations .. the same thing

OpenStudy (anonymous):

Yea , i see .,

OpenStudy (anonymous):

what about the third ? and the last ?

OpenStudy (anonymous):

just calculate and give the answer .. you dont need to show me that you plug 0 in there

OpenStudy (anonymous):

so the 3rd would be : 18(0)^2 + 10(0) + 300 =300 ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

So , now what ? lol

OpenStudy (anonymous):

what about the last ?

OpenStudy (anonymous):

its the same thing as the 3rd .

OpenStudy (anonymous):

good .. so what two equations are good for us so far ?

OpenStudy (anonymous):

1st ?

OpenStudy (anonymous):

1 and 2 .. they are the same so far

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

and both of them passed our first "test" - what is the number of fish in week number 0

OpenStudy (anonymous):

350 ?

OpenStudy (anonymous):

it wasnt a question .. i just said

OpenStudy (anonymous):

ohh lol .

OpenStudy (anonymous):

i feel slow .

OpenStudy (anonymous):

so now we left with P(x) = 13x^2 – 10x + 350; 917 fish P(x) = 13x^2 – 10x + 350; 1,102 fish the difference between those two answers is " Use the model to estimate the number of fish at the lake on week 8." now we have to plug 8 in the equation and see which of them correct is in week number 8 we have 917 fish or 1,102 fish

OpenStudy (anonymous):

So , it will be : 13(8)^2-10(8)+350 =1102

OpenStudy (anonymous):

its the 2nd one (;

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

i really hope you understand what we did

OpenStudy (anonymous):

Thank you , couldn't have done it without you thoughh ! (;

OpenStudy (anonymous):

Yea I did . I understand it now , promise !

OpenStudy (anonymous):

sometimes the process is long and we forget what we are actually looking for

OpenStudy (anonymous):

the first step was eliminating some of the equations using the given data then the second step was checking the validity of the data given in the answer itself

OpenStudy (anonymous):

Yep . . . I see .